# Efficient Offline Reinforcement Learning: The Critic is Critical

### v1

@@ 4.1 Outline Procedure
We now explain our pre-training procedure for the standard return maximising case. In this setting we first compute the discounted return-to-go $R$ from each state-action pair until the end of the trajectory for all timesteps in the dataset. We can then pre-train the actor with behavior cloning and the critic ($Q$-network) with the pre-computed discounted return-to-go, both using supervised mean-squared error (MSE) minimisation (or cross-entropy for discrete actions). We note that for more complex datasets the behavior policy and returns may be asymmetric or multi-modal, in which case the implicit Gaussianity assumption for this loss may be limiting and other pre-training losses may be required, but find MSE is sufficient for most standard environments. This procedure provides initial actor and critic networks consistent with the behavior data. Finally, a suitable off-policy reinforcement learning algorithm (utilizing a temporal difference loss) can be applied to these pre-trained actor and critic networks to efficiently increase the policy return. Our pre-training procedure is outlined in pseudocode in Algorithm 1. As we will see in Section 5, this increase in training efficiency more than makes up for the time and computational expense associated with the pre-training.
4.2 Bias-Variance and Optimism-Pessimism Tradeoffs
Under the behavior policy, this discounted return to go provides a Monte-Carlo (MC) sample of the expectation that the critic or $Q$-network aims to predict. In the case of deterministic environments and a single deterministic behavior policy, this Monte Carlo sample will equal the expectation exactly. For stochastic behavior policies and environments, this Monte Carlo sample may become high variance, which can lead to performance drops after pre-training. In this case it would be possible to use an $n$-step or $\lambda$-return to reduce this variance at the cost of bias introduced by bootstrapping (a well known case of the bias-variance tradeoff) (Sutton et al.,, 2018). However, since computing the $\lambda$-return would require inferring the current critic value for every downstream state in the trajectory, a more practical way of controlling this tradeoff is simply to compute the $TD(0)$ return and combine it with the MC return using a tradeoff parameter $\lambda\in[0,1]$:
@@ 4.2 Bias-Variance and Optimism-Pessimism Tradeoffs
For large offline datasets with high environment coverage where the greatest training efficiency gains are possible, there should be sufficient Monte Carlo samples to reduce this variance to a manageable level for pre-training, so $\lambda$ can be small in order to utilize more information from the offline data (i.e. the full return to go). However, for smaller datasets capturing stochastic policies and environments, larger $\lambda$ may be beneficial. We investigate the empirical effect of varying $\lambda$ in Appendix E.
Additionally, by pre-training on sampled returns with a symmetric error, the critic is equally likely to under- or over-estimate the values of out-of-distribution actions when optimizing the policy after pre-training, even in the deterministic case where the returns are exact. This overestimation can lead to policy collapse as discussed in Section 2. Therefore it can be helpful to add some value regularization $\mathcal{R}(Q(s,a))$ during pre-training, such as that introduced in CQL (Kumar et al.,, 2020) to effectively lower-bound the $Q$ function. We will see in Section 6 that including some value regularization can be beneficial when the offline data is limited. This is reflected in Algorithm 1.
@@
4.3 Generalisation to Maximum Entropy Off-Policy RL Algorithms
The maximum entropy RL framework, and particularly Soft Actor-Critic (Haarnoja et al.,, 2019) along with offline variants such as SAC-N and EDAC (An et al.,, 2021), have recently become popular for their improved robustness and sample efficiency relative to the ‘hard’ return maximisation considered above. This ‘soft’ RL framework involves maximising the expected return alongside the entropy of the policy, balanced by a temperature parameter $\alpha$ (Ziebart et al.,, 2010):
$Q(s_{t},a_{t})=\mathbb{E}_{\tau\sim\pi,\mathcal{T}}\left[\sum_{n=0}^{\infty}\gamma^{n}r_{t+n}+\alpha\sum_{n=1}^{\infty}\gamma^{n}\mathcal{H}(\pi(\cdot|s_{t+n}))\Big{|}s_{t},a_{t}\right]$ (19)
This provides an approximate behavior policy with which we can compute Monte Carlo samples of Equation 19 as soft returns-to-go, which can be used to augment the offline dataset as in Section 4.1. These soft returns-to-go can then be used as the targets to pre-train the critic to achieve consistency with the soft behavior cloned policy, and providing a springboard initialization for a soft off-policy RL algorithm to efficiently improve the policy. The full pseudocode for this procedure is outlined in Algorithm 2, and a further discussion of the rational for this procedure is included in Appendix B.
5 Initial Experiments on D4RL MuJoCo
@@ 5.1 Implementation Details
We begin our investigation into the benefits of pre-training off-policy RL algorithms by considering the D4RL MuJoCo benchmark (Fu et al.,, 2021). We utilize the standard HalfCheetah, Hopper and Walker2d environments and the medium dataset (a suboptimal policy with approximately $1/3$ of the performance of the expert) of 1M transitions, since this provides a meaningful behavior policy to learn from at initialization, with room for improvement with off-policy reinforcement learning. We also consider the medium-replay and full-replay datasets with an identical approach in Appendix F.
@@
For our implementations we utilize the Clean Offline Reinforcement Learning codebase (CORL) (Tarasov et al.,, 2022), that provides algorithm implementations benchmarked to match published performance measures. During our investigation into improving the efficiency of off-policy reinforcement learning algorithms, we found that introducing LayerNorm (Ba et al.,, 2016) into both the actor and critic networks significantly improved training efficiency and stability, independently verifying the findings of Ball et al., (2023), as demonstrated in Figure 2. Therefore for all results demonstrated in this paper, we use the benchmarked implementations and default hyperparameters found in the CORL codebase, with the new addition of a LayerNorm after every linear layer except the final one for each network. We investigate the effect of LayerNorm in more detail in Appendix C.
@@
We use TD3+BC (Fujimoto and Gu,, 2021) as a ‘hard’ off-policy algorithm, and EDAC (An et al.,, 2021) as a ‘soft’ entropy-regularized algorithm to apply after pre-training. For EDAC, we include the auxiliary ensemble diversification loss as value regularization during pre-training, to prevent the collapse of the ensemble. We pre-train until convergence in each case, which can be determined by monitoring the convex supervised loss functions. We find this corresponds to just 10-50k updates and is a small proportion of the total updates required for offline RL convergence. The online performance as a function of offline updates, including pre-training, is demonstrated below in Figure 2.
5.2 Results and Analaysis
Figure 2: Performance mean $\pm$ standard deviation at each training step over 3 independent seeds. Supervised pre-training before offline reinforcement learning is more efficient than offline reinforcement learning from scratch. Surprisingly, performance is also more stable long after pre-training.
We find that pre-training as described in Algorithms 1, 2 leads to much more efficient training, both for TD3+BC (a hard RL algorithm with actor regularization) and for EDAC (a soft RL algorithm with critic regularization), even when taking the cost of pre-training into account. In particular, we find that on the more difficult Hopper and Walker2d environments, the inclusion of pre-training generally reaches expert-level performance in less than $\mathbf{1/2}$ of the training steps and computation time required without pre-training (and in less than $1/5$ of the training steps and computation time of the currently used implementations without LayerNorm or pre-training). Surprisingly, we notice that in many cases the final performance is also more stable, which we discuss further in Appendix D.
However, in some cases, we find that the performance drops after critic pre-training. Fundamentally, this arises because at the end of pre-training we change the objective of the actor from imitation learning (predicting the action that would have been chosen in the dataset), to off-policy reinforcement learning (predicting the action that will maximise the critic prediction). If the values predicted by the critic are sufficiently accurate for the behavior policy as a result of our proposed critic pre-training, then the performance should smoothly improve, but if values are inaccurate due to sample variance then the performance may drop and take some time for Bellman updates to reduce this variance. The dynamics of these environments are deterministic, but the D4RL datasets are collected with a stochastic policy (SAC, Haarnoja et al., (2018)) and from stochastic initial states, which leads to variance in the MC value estimates. Since the action spaces are larger for HalfCheetah and Walker2d than for Hopper (6 dimensional rather than 3), there will be more variance in the policy at each timestep and therefore also in the return. Furthermore, since all but one of the medium-level Hopper trajectories end in termination, this increases the return signal-to-noise ratio for Hopper relative to the HalfCheetah and Walker2d environments which end with timeouts. To mitigate the performance drop off in these environments, we incorporate a small amount of TD target using Equation 17, with $\lambda=0.1$ for HalfCheetah, and Walker2d for EDAC. We investigate the inclusion of $\lambda$ in Appendix E.
As an ablation, we also investigate pre-training the actor alone (BC only pre-training) in Appendix D. We find that performance quickly deteriorates after pre-training due to the inconsistent random critic values, verifying our hypothesis on the performance drop, and the findings of Orsini et al., (2021).
6 Extension to Data-Limited Adroit Environments
We now consider the Adroit tasks (Rajeswaran et al.,, 2018). These are more complex and realistic environments that require controlling a 24-DoF robotic hand to perform tasks such as aligning a pen, hammering a nail, opening a door, or relocating a ball. The human datasets provided for these environments are very limited, consisting of just 25 trajectories of human demonstrations. The cloned datasets augment these trajectories with a behavior cloned policy to get a 50-50 mixture. These demonstrations can be improved upon by acting more efficiently, but are much better than random behavior, so the optimal policy and values should be close to those of the behavior data.
6.1 Motivation for Actor and Critic Regularization With Pre-Training
In such data-limited settings however, off-policy algorithms often suffer from policy collapse as discussed in Section 4.1, since the actor or critic erroneously extrapolate out-of-distribution (OOD) of the offline data. Indeed, when we applied our pre-training approach from Section 5 to TD3+BC and EDAC on the Adroit environments, we found that the human-level pre-training performance often rapidly collapsed after pre-training. As we saw in Section 5.2 and Appendix D, consistency of both actor and critic are crucial for performance improvement. However, most offline RL methods only apply regularization to only one of either the actor or the critic (as also noted by Tarasov et al., (2023)). Motivated by this insight, we introduce two new algorithms that incorporate regularization on both the actor and the critic to enable smooth performance improvement after pre-training in data-limited domains. First, we introduce TD3+BC+CQL, which combines the existing behavior cloning on the actor, with additional CQL regularization on the critic, to penalise large OOD Q-values. Second, we introduce EDAC+BC, which combines the existing uncertainty-based regularization on the critic, with additional behavior cloning on the actor, to penalise OOD actions.
6.2 Implementation Details
For all baselines we use their benchmarked CORL implementations (Tarasov et al.,, 2022) and previously published hyperparameters where possible. Crucially, for the TD3+BC and EDAC baselines, we use the same regularization as TD3+BC+CQL and EDAC+BC for the existing regularization components. We train all algorithms for 300k updates (corresponding to up to 4 hours training time on our RTX2080 GPUs). For our novel algorithms, we pre-train for 200k steps, to provide ample time for supervised convergence. To fairly measure performance we then average online performance evaluated every 10000 offline updates between 200k and 300k updates, over 4 independent seeds. Full details of our hyperparameters and procedure are provided in Appendicies G and H.

### v2

@@ 5.1 Outline Procedure
We now explain our pre-training procedure for standard return maximizing algorithms. In this case, we first compute the discounted return-to-go $R$ from each state-action pair until the end of the trajectory for all timesteps in the dataset. We can then pre-train the actor with behavior cloning and the critic ($Q$-network) with the pre-computed discounted return-to-go $R$, both using supervised mean-squared error (MSE) minimization (or alternatively cross-entropy for discretized actions/values). This procedure provides consistent initial actor and critic networks. Subsequently, a suitable off-policy RL algorithm (using a temporal-difference loss) can be applied to these pre-trained actor and critic networks to efficiently increase the policy return. Our pre-training procedure is outlined in pseudocode in Algorithm 1. As we will see in Section 6, this increase in training efficiency more than makes up for the time and computational expense associated with the supervised pre-training.
5.2 Bias-Variance and Optimism-Pessimism Trade-offs
Under the behavior policy, this discounted return to go provides a Monte-Carlo (MC) sample of the expectation that the critic or $Q$-network aims to predict. In the case of deterministic environments and a single deterministic behavior policy, this Monte Carlo sample will equal the expectation exactly. For stochastic behavior policies and environments, this Monte Carlo sample may become high variance for finite sample datasets, which can lead to performance drops after pre-training. In this case it would be possible to use an $n$-step or $\lambda$-return to reduce this variance at the cost of bias introduced by bootstrapping (a well known case of the bias-variance trade-off) (sutton_reinforcement_2018). However, since computing the $\lambda$-return would require inferring the current critic value for every downstream state in the trajectory, an alternative way of controlling this trade-off is simply to compute the $TD(0)$ return and combine it with the MC return using a trade-off parameter $\lambda\in[0,1]$:
@@ 5.2 Bias-Variance and Optimism-Pessimism Trade-offs
For large offline datasets with good environment coverage, where the greatest training efficiency gains are possible, there should be sufficient MC samples to reduce this variance to a manageable level for pre-training, so $\lambda$ can be small in order to utilize more information from the offline data (i.e. the full return-to-go). However, for smaller datasets capturing stochastic policies and environments, larger $\lambda$ may be beneficial. We investigate the empirical effect of varying $\lambda$ in our experiments.
Additionally, by pre-training on sampled returns with a symmetric error, the critic is equally likely to under- or over-estimate the values of out-of-distribution actions when optimizing the policy after pre-training, even in the deterministic case where the returns are exact. This overestimation can lead to policy collapse as discussed in Section 2. Therefore it can be helpful to add some value regularization $\mathcal{R}(Q(s,a))$ during pre-training, such as that introduced in CQL (kumar_conservative_2020) to effectively lower-bound the $Q$ function. We find in our experiments that including some value regularization can be beneficial when the offline data is limited. This is reflected in Algorithm 1.
5.3 Generalization to Maximum Entropy Off-Policy RL Algorithms
The maximum entropy RL framework, and particularly Soft Actor-Critic (haarnoja_soft_2019) along with offline variants such as SAC-N and EDAC (an_uncertainty-based_2021), have recently become popular for their improved robustness and sample efficiency relative to the ‘hard’ return maximization considered above. This ‘soft’ RL framework involves maximizing the expected return alongside the entropy of the policy, balanced by a temperature parameter $\alpha$ (ziebart_modeling_2010):
$Q(s_{t},a_{t})=\mathbb{E}_{\tau\sim\pi,\mathcal{T}}\left[\sum_{n=0}^{\infty}\gamma^{n}r_{t+n}+\alpha\sum_{n=1}^{\infty}\gamma^{n}\mathcal{H}(\pi(\cdot|s_{t+n}))\Big|s_{t},a_{t}\right].$ (19)
This provides an approximate behavior policy with which we can compute Monte Carlo samples of Equation 19 as soft returns-to-go, which can be used to augment the offline dataset as in Section 5.1. These soft returns-to-go can then be used as the targets to pre-train the critic to achieve consistency with the soft behavior cloned policy, and providing a springboard initialization for a soft off-policy RL algorithm to efficiently improve the policy. The full pseudocode for this procedure is outlined in Algorithm 2, and a further discussion of the rational for this procedure is included in Appendix A.
@@ 5.3 Generalization to Maximum Entropy Off-Policy RL Algorithms
6 Experiments on D4RL MuJoCo
@@ 6.1 Implementation Details
To experimentally validate the hypothesized benefits of pre-training off-policy RL algorithms, we consider the standard D4RL MuJoCo benchmark (fu_d4rl_2021). We use the standard HalfCheetah, Hopper and Walker2d environments and the medium dataset (a suboptimal policy with approximately $1/3$ of the performance of the expert) of 1M transitions, since this provides a meaningful behavior policy to learn from at initialization, with room for improvement with off-policy reinforcement learning. We also investigate the effect of pre-training on the medium-replay and full-replay datasets in Appendix E, which also serve as sub-optimal but non-random behavior datasets.
@@
For our implementations we utilize the Clean Offline Reinforcement Learning codebase (CORL) (tarasov2022corl), that provides algorithm implementations benchmarked to match published performance measures. During our experiments, we found that introducing LayerNorm (ba_layer_2016) into both the actor and critic networks significantly improved training efficiency and stability, verifying the findings of ball_efficient_2023. Therefore, for all results demonstrated in this paper, we use the benchmarked implementations from the CORL codebase, with the new addition of a LayerNorm after every linear layer except the final one for each network. We investigate the effect of LayerNorm in more detail in Appendix B.
@@
We use TD3+BC (fujimoto_minimalist_2021) as a ‘hard’ off-policy algorithm, and EDAC (an_uncertainty-based_2021) as a ‘soft’ entropy-regularized algorithm to apply after pre-training. For EDAC, we include the auxiliary ensemble diversification loss as value regularization during pre-training, to prevent the collapse of the ensemble. We pre-train until convergence in each case, which can be determined by monitoring the convex supervised loss functions. Crucially, this corresponds to just 10-50k updates and is a small proportion of the total updates required for offline RL convergence. The online performance as a function of offline updates, including pre-training, is demonstrated below in Figure 2.
@@ 6.2 Results and Analysis
Figure 2: Performance means with shaded 95% confidence intervals across 3 independent seeds. Supervised pre-training before offline reinforcement learning is more efficient than offline reinforcement learning from scratch. Surprisingly, performance is often more stable long after pre-training.
We find that pre-training as described in Algorithms 1, 2 leads to much more efficient training, both for TD3+BC (a hard RL algorithm with actor regularization) and for EDAC (a soft RL algorithm with critic regularization), even when taking the cost of pre-training into account. In particular, we find that on the more difficult Hopper and Walker2d environments, the inclusion of pre-training generally reaches asymptotic performance in less than $\mathbf{1/2}$ of the training steps and computation time required without pre-training. Surprisingly, we notice that in many cases the final performance is also more stable, which we investigate further with ablations in Appendix C. We find similar benefits from pre-training on the medium-replay and full-replay datasets in Appendix E.
In some cases, we found that the performance drops after critic pre-training. Fundamentally, this arises because at the end of pre-training we change the objective of the actor from imitation learning (predicting the action that would have been chosen in the dataset), to off-policy reinforcement learning (predicting the action that will maximize the critic prediction). If the values predicted by the critic are sufficiently accurate around the behavior policy after critic pre-training, the performance should smoothly improve, but if values are inaccurate due to variance then the performance may drop and take some time for Bellman updates to reduce this variance. In particular, the dynamics of these environments are deterministic, but the D4RL datasets are collected with a stochastic policy (SAC, haarnoja_soft_2018) and from stochastic initial states, which leads to variance in the MC value estimates. Since the action spaces are larger for HalfCheetah and Walker2d than for Hopper (6 dimensional rather than 3), there will be more variance in the policy at each timestep, since the target policy entropy for SAC is the dimension of the actor space $\dim(A)$. This causes greater variance in the return. Furthermore, since all but one of the medium-level Hopper trajectories end in termination, this increases the return signal-to-noise ratio for Hopper relative to the HalfCheetah and Walker2d environments which end with timeouts. To mitigate the performance drop off in these environments, we incorporate a small amount of TD target using Equation 17, with $\lambda=0.1$ for HalfCheetah, and Walker2d for EDAC. We investigate the effect of $\lambda$ fully in Appendix D.
To further validate the benefit of pre-training, we additionally investigate the effect of pre-training on the Adroit environments (rajeswaran_learning_2018) in Appendix G. These are more complex and realistic tasks that require controlling a 24-DoF robotic hand to perform tasks such as aligning a pen, hammering a nail, opening a door, or relocating a ball. These form suitable environments for our approach, since the provided human demonstrations are much better than random behavior (and are therefore worth imitating) but can be improved upon by acting more efficiently, so it is possible to improve on this imitation policy. However, the human datasets provided for these environments are very limited, consisting of just 25 trajectories of human demonstrations per task. This means that additional regularization towards the behavior policy is necessary to prevent performance collapse after pre-training for the standard off-policy algorithms we consider. With this additional regularization in place, these experiments provide further evidence for the benefits of pre-training policies and values when learning a policy from effective but suboptimal human demonstrations.
7 Conclusion
We have demonstrated that pre-training off-policy algorithms to imitate the behavior policy and corresponding values to obtain a consistent actor and critic before attempting improvement can increase the computational efficiency and stability of subsequent off-policy reinforcement learning. In particular, we have shown that convergence should theoretically improve logarithmically as the pre-trained initialization improves (as measured by the infinite norm of the behavior policy values with respect to the optimal values). We have demonstrated experimentally that these efficiency improvements hold in practice for modern actor-critic algorithms, and can more than halve the computation time required for these algorithms to converge on standard offline reinforcement learning environments. Our proposed additional pre-training phase can be widely applied to current and future off-policy reinforcement learning algorithms and requires minimal algorithmic modification. We hope that our research contributes to bridging the gap between theoretically-tractable tabular algorithms and empirically driven deep reinforcement learning, while improving the efficiency and stability of offline reinforcement learning as scale continues to increase.
Broader Impact Statement
Our work considers making offline reinforcement learning more computationally efficient and stable. The authors are not aware of any additional ethical concerns arising from our contributions that are not present in existing methods.
Acknowledgments and Disclosure of Funding
Appendix A Rational for Separation of Actor and Critic Pre-training for Entropy-Regularized Reinforcement Learning
In Section 5.3 and Algorithm 2 of the main text, we propose separating the pre-training of the policy and value network into two separate phases for entropy-regularized reinforcement learning algorithms. By first pre-training the policy with soft behavior cloning, an approximate behavior policy can be learned which then enables approximate behavior entropy bonuses to be included in the subsequent pre-training of the critic. However, an alternative approach could be to pre-train the policy and critic in parallel, as in the pure return maximization framework. In theory this would require updating the returns-to-go for each policy update to incorporate the changing entropy bonuses as the variance of the policy is updated. Since this requires a complete forward pass of the policy for all subsequent states in the trajectories of the states sampled for an update, this makes the pre-training significantly more computationally expensive.
Another potential approach could be to only train the mean of the standard Gaussian policy to match the behavior policy, and keep the variance constant such that all entropy bonuses could be calculated at initialization and would be unaffected by policy pre-training. However, the standard tanh squashing applied to the policy to keep the sampled action within the environment action bounds leads to a changing entropy of the resulting policy, even with the Gaussian variance kept constant.
A third approach could be to separate the policy and value pre-training, but continue the policy behavior cloning during value pre-training. While this is possible, for the entropy bonuses to be sufficiently accurate, the policy behavior cloning must have effectively converged before value pre-training, meaning continuing to pre-train the policy in parallel with the critic has little effect and requires unnecessary computation time.
A final approach could be to compute the soft returns-to-go based on the initialization policy, and then only pre-train the values (no behavior cloning). While this approach was successful and led to training efficiency gains, the rapid updating of the actor at the beginning of training (and particularly the rapidly changing policy entropy) quickly leads to inconsistent values, so we found that the investment in pre-training the policy with soft behaviour cloning first was worth the computational time in most cases.
Appendix B Investigation Into Affect of LayerNorm
Figure 3: Investigation into the affect of adding LayerNorm to both the actor and critic networks for TD3+BC on HalfCheetah and Hopper-medium. All lines show mean and standard deviation in normalized return at each timestep over 3 seeds.
We investigate the effect of the addition of LayerNorm (ba_layer_2016) to the actor and critic networks individually for TD3+BC on the HalfCheetah-medium and Hopper-medium datasets. The original author implementation of TD3+BC (along with that of SAC-N, EDAC and many other RL algorithms (tarasov2022corl)) does not include any form of normalization, and is shown in green. We consider the addition of LayerNorm after every linear layer in the network (before activation) except the final linear layer. We find that adding LayerNorms to the critic network leads to significant improvement in training efficiency and stability. This independently verifies the findings of ball_efficient_2023, who hypothesize that this occurs because the normalization prevents severe value extrapolation for out-of-distribution actions, leading to overestimation error. Surprisingly, the addition of LayerNorm to the actor alone leads to worse efficiency and stability than no LayerNorm for the HalfCheetah environment. However, the addition of LayerNorm to both the actor and the critic leads to greater training efficiency and stability than the default, for both environments.
We find that these insights generally hold across investigated environments and datasets, as shown in Figure 4 below. We therefore apply LayerNorm to both the actor and the critic for all experiments in this paper towards our goal of improving training efficiency. We expect the addition of LayerNorm to be universally used for off-policy reinforcement learning algorithms going forwards.
Figure 4: The effect of adding LayerNorm to both actor+critic across environments. All lines show mean and standard deviation in normalized return at each timestep over 3 seeds.
Appendix C Ablations of Actor and Critic Pre-Training
Figure 5: Ablation of actor and critic pre-training. All lines show mean and standard deviation in normalized return at each timestep over 3 seeds.
We see that when only pre-training the actor with behavior cloning, the initial performance matches our proposed actor and critic pre-training, but quickly declines after pre-training due to the randomly initialized critic, and is no better than training from scratch. Pre-training the critic is much more important for efficient performance improvement, and is better than no pre-training, but performance starts from random so is not as effective as combined actor and critic pre-training. This demonstrates that having a consistent actor and critic is important for improved training efficiency and stability.
Surprisingly, we also notice that in many cases the asymptotic performance with combined actor and critic pre-training is more stable, even after hundreds of thousands of updates after pre-training. As analyzed in fujimoto_why_2022 and chen_information-theoretic_2019, for finite data regimes such as the offline setting, the Bellman equation can be satisfied by infinitely many suboptimal solutions. We hypothesize that this additional stability could be occurring because the initial pre-training to obtain a consistent actor and critic reduces the subset of possible solutions to those with lower value error when subsequently minimizing the Bellman error on the finite offline dataset. However, since this benefit is auxiliary to our central focus of improving efficiency, we leave investigation of this effect to further work.
Appendix D Investigation into Empirical Bias Variance Tradeoff
Figure 6: Investigation into the empirical bias-variance tradeoff by varying $\lambda$ defined in Equation 17. All lines show mean and standard deviation in normalized return at each timestep over 3 seeds.
We investigate the bias-variance trade-off described in Section 5.2 in practice by empirically varying $\lambda$ across algorithms and environments. We find that while all values of $\lambda\in[0,1]$ provide efficiency benefits over no pre-training, the effect of $\lambda=0$ (corresponding to fully supervised value-error pre-training) is often quite different, likely because even for $\lambda=0.1$ the temporal difference component to the loss can have significant impact on the training dynamics, and the pre-training duration is not long enough for this bootstrapping loss to reach consistency. However, including some temporal difference component ($\lambda>0$) can help to smooth the transition between imitation learning and off-policy reinforcement learning, as we see particularly evidently for the HalfCheetah environment.
As discussed in the main text in Section 6, if the values predicted by the critic are sufficiently accurate around the behavior policy after critic pre-training, then the performance should smoothly improve, but if values are inaccurate due to variance or extrapolation error, then the performance may drop and take some time for Bellman updates to reduce this variance. In particular, the dynamics of these environments are deterministic, but the D4RL datasets are collected with a stochastic policy (SAC, haarnoja_soft_2018) and from stochastic initial states, which leads to variance in the MC value estimates. Since the action spaces are larger for HalfCheetah and Walker2d than for Hopper (6 dimensional rather than 3), there will be more variance in the policy at each timestep, since the target policy entropy for SAC is the dimension of the actor space $\dim(A)$. This causes greater variance in the return. Furthermore, since all but one of the medium-level Hopper trajectories end in termination, this increases the return signal-to-variance ratio for Hopper relative to the HalfCheetah and Walker2d environments, which end with timeouts. More generally, we expect that larger scale datasets with broad coverage of the relevant state-action space to have less variance in the MC return targets and therefore have less need for a TD component ($\lambda>0$).
Appendix E MuJoCo Medium-Replay and Full-Replay Dataset Experiments
We apply our pre-training approach proposed in Section 5 with an identical experimental implementation to that described in Section 6.1 to the medium-replay and full-replay datasets, shown below in Figures 7 and 8 respectively.
E.1 Medium-replay
The medium-replay dataset consists of 1M transitions from the replay buffer of an agent trained from random to medium performance. The effect of combined actor and critic pre-training on these datasets is shown below.
We generally see similar efficiency and stability gains to those observed for the medium datasets in Figure 2. We also notice that the performance of Walker2d for EDAC is much cleaner, likely due to greater diversity of data helping to stabilize performance.
E.2 Full-replay
The full-replay dataset consists of 1M transitions from the replay buffer of an agent trained from random to expert performance. The effect of combined actor and critic pre-training on these datasets is shown below.
As with the medium-replay datasets we see similar efficiency and stability gains to those observed for the medium datasets in Figure 2 (with the exception of the HalfCheetah environments). In some cases, the efficiency gains may be further improved by optimization of the pre-training duration and temporal-difference component $\lambda$ from those that were originally selected for the medium datasets.
Appendix F Investigation of Alternative Pre-training Method used by $Q$-Transformer
As mentioned in Section 2, $Q$-Transformer (chebotar_q-transformer_2023) recently claimed improved training efficiency by incorporating MC values into the Bellman target via a max operation of the form $\max(MC,Q)$ (as originally proposed by wilcox_monte_2022). We compare this objective to our combined actor and critic pre-training objective in Figure 9 below.
Figure 9: Comparison of the $Q$-Transformer training objective incorporating MC values via a $\max$ operation with our combined actor and critic pre-training. Plots show mean and standard deviation at each timestep for 3 independent runs.
We find that our proposed combined actor and critic pre-training generally performs better than the $Q$-Transformer objective. Given the claimed performance benefit of this objective in $Q$-Transformer, we hypothesize that this relatively poor performance may be due to the variance in the MC values that is observed with these smaller scale datasets. In particular, in this smaller-scale offline setting, some MC values may be high variance which can prevent a clean transition from MC to TD ($Q$) values and prevent the temporal-difference loss from being properly optimized. We note that an effective hybrid approach might be to use this objective for pre-training, and end the pre-training when a certain threshold on the proportion of MC targets being selected by the max operation is reached, alleviating the ongoing impact of the MC values and the need for a pre-defined pre-training period. However, we leave a deeper investigation of this alternative objective to future work.
Appendix G Extension to Data-Limited Adroit Environments
In addition to our main MuJoCo experiments in Section 6, we consider the Adroit environments (rajeswaran_learning_2018), which are more complex and realistic tasks that require controlling a 24-DoF robotic hand to perform tasks such as aligning a pen, hammering a nail, opening a door, or relocating a ball. These form suitable environments for our approach, since the provided human demonstrations are much better than random behavior (and are therefore worth imitating initially) but can be improved upon by acting more efficiently, so it should be possible to improve on this imitation policy. However, the human datasets provided for these environments are very limited, consisting of just 25 trajectories of human demonstrations per task. The cloned datasets augment these trajectories with a behavior cloned policy to get a 50-50 mixture.
G.1 Motivation for Actor and Critic Regularization With Pre-Training
In such data-limited settings however, off-policy algorithms often suffer from policy collapse as discussed in Section 5.1, since the actor or critic erroneously extrapolate out-of-distribution (OOD) of the offline data. Indeed, when we applied our pre-training approach from Section 6 to TD3+BC and EDAC on the Adroit environments, we found that the human-level pre-training performance often rapidly collapsed after pre-training. As we saw in Section 6.2 and Appendix C, consistency of both actor and critic are crucial for performance improvement. However, most offline RL methods apply regularization to only one of either the actor or the critic (as also noted by tarasov_revisiting_2023). Motivated by this insight, we introduce two new algorithms that incorporate regularization on both the actor and the critic to enable smooth performance improvement after pre-training in data-limited domains. First, we introduce TD3+BC+CQL, which combines the existing behavior cloning on the actor, with additional CQL regularization on the critic, to penalize large OOD Q-values. Second, we introduce EDAC+BC, which combines the existing uncertainty-based regularization on the critic, with additional behavior cloning on the actor, to penalize OOD actions.
G.2 Implementation Details
For all baselines we use their benchmarked CORL implementations (tarasov2022corl) and previously published hyperparameters where possible. Crucially, for the TD3+BC and EDAC baselines, we use the same regularization as TD3+BC+CQL and EDAC+BC for the existing regularization components. We train all algorithms for 300k updates (corresponding to up to 4 hours training time on our RTX2080 GPUs). For our novel algorithms, we pre-train for 200k steps, to provide ample time for supervised convergence. To fairly measure performance we then average online performance evaluated every 10000 offline updates between 200k and 300k updates, over 4 independent seeds. Full details of our procedure and hyperparameters are provided in Appendices G.5 and G.4.